OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue). A. 4/3 B. 2/3 C. 7/9 D. 4/27

OpenStudy (anonymous):

thought we just did this one yes?

OpenStudy (anonymous):

\[\frac{4}{9}\times \frac{3}{9}=\frac{4}{9}\times \frac{1}{3}=...\]

OpenStudy (anonymous):

whats the answer